### v1

@@ Abstract
Planetary surface habitability has so far been, in the main, considered in its entirety. The increasing popularity of 3D modelling studies of (exo)planetary climate has highlighted the need for a new measure of surface habitability. Combining the observed thermal limits of modern Earth-based life with surface water fluxes, we introduce such a measure which can be calculated from the climatological outputs from general circulation model simulations. In particular, we pay attention to not only the thermal bounds of macroscopic complex life, but additionally the limits of microbial and extremophilic life which have been vital to the generation of Earth’s own biosignatures. This new metric is validated on Earth, using ERA5 reanalysis data to predict the distribution of surface habitability which is compared to the observed habitability generated from satellite-derived data of photosynthetic life. The validation against observed habitability is additionally repeated for a selection of other defined metrics of surface habitability, allowing for the first time a comparison of metric performance with respect to Earth-based surface life, finding that future climatologically-based metric definitions should include both conditions for temperature and water/nutrient availability.
@@ 4.3.2 Observed habitability calculation
$H_{O_{T}}$ =$=\begin{cases}\text{complex}&\text{if}\quad\text{NDVI}_{\text{mean}}>0.3,\\ \text{microbial}&\text{if}\quad\text{NDVI}_{\text{mean}}>0.15,\\ \text{limited}&\text{otherwise},\end{cases}$ (5)
@@ 5 Conclusion
We have reviewed a set of climate-based metrics of surface habitability against the distribution of photosynthetic life on Earth, finding that a metric defined using either surface air temperature $T_{s}$, aridity index $AI$, or sea ice concentration $SIC$ alone is not sufficient in capturing the pattern of habitability on Earth. Metrics dependent upon $T_{s}$ and $SIC$ approximately capture the shift to microbial and limited habitability at higher latitudes, but incorrectly label regions of observed limited habitability regions at lower latitudes as habitable, resulting in an overestimation in the fractional habitability $f_{H}$. The metrics dependent upon $AI$ show different patterns of habitability to $T_{s}$ which begin to resemble some of the lower latitude terrestrial regions of limited habitability but exhibit poor predictive skill overall. We find that our newly defined metric, defined using $T_{s}$, precipitation $P$, and evaporation $E$, shows a vast improvement to the other metrics, with the highest accuracy, predictive skill, and (mostly) closest $f_{H}$ to the observed. Globally, our metric has accuracies of 0.67 and 0.70 for microbial and complex habitability respectively, and performs particularly well in the terrestrial domain with respective accuracies of 0.77 and 0.80. Disagreements between our metric and the observed can be attributed to a mixture of uncertainties from the data sources used to calculate observed habitability, seasonal variability, as well as the metric failing to entirely capture the distribution of nutrients across the marine domain. The incorporation of additional climate parameters may serve to better describe the global distribution of nutrients as well as surface water fluxes, however we leave this for future work to investigate, noting the simplicity of the metric in its current form. Finally, we highlight that use of our metric considers that other requirements of habitability, including active carbon cycle and plate tectonics, suitable atmospheric composition, and spectral energy distribution, as satisfied. In this way, we suggest that the metric be used in tandem with other measures of habitability, including but not limited to, carbon cycle models, Habitable Zone and Habitable Zone Lifetimes, Earth Similarity Index, Gaia modelling, or nutrient availability [119, 120, 121, 5, 122].
@@ Acknowledgements
This work was supported by NERC via the London NERC DTP, grant number NE/S007229/1. N.J.M. acknowledges support from a UKRI Future Leaders Fellowship [Grant MR/T040866/1], a Science and Technology Facilities Funding Council Nucleus Award [Grant ST/T000082/1], and the Leverhulme Trust through a research project grant [RPG-2020-82].

### v2

@@ Abstract
Planetary surface habitability has so far been, in the main, considered in its entirety. The increasing popularity of 3D modelling studies of (exo)planetary climate has highlighted the need for a more nuanced understanding of surface habitability. Using satellite-derived data of photosynthetic life to represent the observed surface habitability of modern Earth, we validate a set of climatologically-defined metrics previously used in exoplanetary habitability studies. The comparison finds that the metrics defined by temperature show spatial patterns of habitability distinct to those defined by aridity, with no metric able to completely replicate the observed. We build upon these results to introduce a new metric defined by the observed thermal limits of modern Earth-based life, along with surface water fluxes as an analogue for water and nutrient availability. Furthermore, we pay attention to not only the thermal bounds of macroscopic complex life, but additionally the limits of microbial life which have been vital to the generation of Earth’s own biosignatures, thus expanding considerations of climatic habitability out of a historically binary definition. Repeating the validation for our metric finds a significant improvement in the spatial representation of habitability, laying the groundwork for more accurate assessments of potential life-supporting environments upon other planets.
@@ 4.3.2 Observed habitability calculation
$H^{obs}_{T}$ =$=\begin{cases}\text{complex}&\text{if}\quad\text{NDVI}_{\text{mean}}>0.3,\\ \text{microbial}&\text{if}\quad\text{NDVI}_{\text{mean}}>0.15,\\ \text{limited}&\text{otherwise},\end{cases}$ (5)
@@ 5 Conclusion
We have reviewed a set of climate-based metrics of surface habitability against the distribution of photosynthetic life on Earth, finding that a metric defined using either surface air temperature $T_{s}$, aridity index $AI$, or sea ice concentration $SIC$ alone is not sufficient in capturing the pattern of habitability on Earth. Metrics dependent upon $T_{s}$ and $SIC$ approximately capture the shift to microbial and limited habitability at higher latitudes, but incorrectly label regions of observed limited habitability regions at lower latitudes as habitable, resulting in an overestimation in the fractional habitability $f_{H}$. The metrics dependent upon $AI$ show different patterns of habitability to $T_{s}$ which begin to resemble some of the lower latitude terrestrial regions of limited habitability but exhibit poor predictive skill overall. We find that our newly defined metric, defined using $T_{s}$, precipitation $P$, and evaporation $E$, shows a vast improvement to the other metrics, with the highest accuracy, predictive skill, and (mostly) closest $f_{H}$ to the observed. Globally, our metric has accuracies of 0.67 and 0.70 for microbial and complex habitability respectively, and performs particularly well in the terrestrial domain with respective accuracies of 0.77 and 0.80. Disagreements between our metric and the observed can be attributed to a mixture of uncertainties from the data sources used to calculate observed habitability, seasonal variability, as well as the metric failing to entirely capture the distribution of nutrients across the marine domain. The incorporation of additional climate parameters may serve to better describe the global distribution of nutrients as well as surface water fluxes, however we leave this for future work to investigate, noting the simplicity of the metric in its current form. Finally, we highlight that use of our metric considers that other requirements of habitability, including active carbon cycle and plate tectonics, suitable atmospheric composition, and spectral energy distribution, as satisfied [30, 7]. In this way, we suggest that the metric be used in tandem with other measures of habitability, including but not limited to, carbon cycle models, Habitable Zone and Habitable Zone Lifetimes, Gaia modelling, or nutrient availability [119, 120, 121, 5].
@@ Acknowledgements
The authors thank Ian Crawford for providing constructive feedback on preliminary drafts of this work. H.L.W. acknowledges support from NERC via the London NERC DTP [Grant NE/S007229/1]. N.J.M. acknowledges support from a UKRI Future Leaders Fellowship [Grant MR/T040866/1], a Science and Technology Facilities Funding Council Nucleus Award [Grant ST/T000082/1], and the Leverhulme Trust through a research project grant [RPG-2020-82].